Question 3 (10 marks) Atlas Pty Ltd needs $50 million for its next growth phase. It plans to raise the money by an Initial Public Offering (IPO) of shares and has provided the following information to the market The first dividend of $2.50 will be paid next vear . After that dividend, dividends are expected to grow annually by 4% per annum, in perpetuity The underwriters will charge a 7 per cent spread. Assume shareholders require a return of 16% per annum (effective annual rate). a) Calculate the intrinsic value of a share in Atlas (3 marks) If the offer price to the public is $16.00 per share, calculate the number of shares that Atlas will need to issue to achieve its goal (of $50million, net of the underwriters fee) b) (3 marks)
Answer #1

A) By dividend discount model for constant growing dividends, current intrinsic value of share is present value of future dividends expected from next year onwards.

Hence, V0 = D1/(r – g)

D1 = $2.50

R = 16%, g = 4%

V0 = 2.50/(12% - 4%) = $20.83. Answer

B) Offer Price = $16, Underwriting spread = 7%

This implies, the firm will receive $16 * (1 -7%) = $14.88 per share

If the company wants to raise $50 mil, and they would be receiving $14.88, the number of shares required to be issued = 3,360,215 shares. Answer


answered by: ANURANJAN SARSAM